# Carefree multiple testing with e-processes

### v1

@@ Abstract
E-processes enable hypothesis testing with ongoing data collection while maintaining Type I error control. However, when testing multiple hypotheses simultaneously, current $e$-value based multiple testing methods such as e-BH are not invariant to the order in which data are gathered for the different $e$-processes. This can lead to undesirable situations, e.g., where a hypothesis rejected at time $t$ is no longer rejected at time $t+1$ after choosing to gather more data for one or more $e$ processes unrelated to that hypothesis. We argue that multiple testing methods should always work with suprema of $e$-processes. We provide an example to illustrate that e-BH does not control this FDR at level $\alpha$ when applied to suprema of $e$-processes. We show that adjusters can be used to ensure FDR-sup control with e-BH under arbitrary dependence.
@@ 1 Introduction
Unfortunately, data gathering under the guarantee (2) is not carefree. Let us illustrate this with an example. For $K=2$ $e$-values, e-BH rejects any hypothesis for which the $e$-value exceeds $2/\alpha$, and both hypotheses if both $e$-values exceed $1/\alpha$. Now suppose that at time $t$ we have $1/\alpha\leq E_{t}^{1}<2/\alpha$ and $E_{t}^{2}<1/\alpha$. Gathering more data for both $e$-processes, however, the situation could reverse, and $E_{t+1}^{1}<1/\alpha$ and $1/\alpha\leq E_{t+1}^{2}<2/\alpha$. Now, the researcher regrets gathering more data for $E^{1}$, since (2) does not support any rejection now, but would have supported rejection of both null hypotheses if the researcher would have stopped $E^{1}$, taking $E^{1}_{t+1}=E^{1}_{t}$. The lack of the carefree property hugely complicates experimental design.
For carefree data gathering, the researcher must always be allowed to revert any $e$-process back to an earlier time point. The optimal reversion is to the maximum so far, so a carefree procedure should guarantee
@@ 2.1 The e-BH procedure
Formally, a non-negative process $E$ is called an $e$-process when $\mathbb{E}[E_{\tau}]\leq 1$ for any stopping time $\tau$, where the expectation is taken under any distribution in the null hypothesis. For each $k\in\mathcal{K}$, let $H_{k}$ be associated with an e-process $(E^{k}_{t})_{t\geq 0}$, and let $\tau_{1},\tau_{2},\ldots,\tau_{K}$ be their arbitrary stopping times. Let $S_{k}=E^{k}_{\tau_{k}}$ be the stopped $e$-processes.
@@
The stopped e-BH procedure applied to arbitrarily dependent e-processes has FDR at most $K_{0}\alpha/K$.
@@ 2.3 Related work
Online multiple testing has been considered both classically with $p$-values [Fischer et al., 2024] and $e$-values [Fischer and Ramdas, 2024a, b, Xu and Ramdas, 2023], but online is here to be understood in the dimension of the hypothesis space: time points represent the addition of an hypothesis. The only paper to our knowledge that considers multiple testing with $e$-values where data can be added sequentially, is [Xu and Ramdas, 2023], where they introduce the doubly sequential framework. For $p$-values there is work on interim-analyses and group-sequential multiple testing, e.g. [Sarkar et al., 2019] on FDR; but to the best of our knowledge there is no work on any-time valid methods.
3 FDR with the running maximum of an e-process
@@ 3.1 Proof of Proposition 1
$M_{t}^{1}\geq\frac{2}{\alpha},\quad M_{t}^{2}\geq\frac{2}{\alpha},\quad\text{or both }M_{t}^{1},M_{t}^{2}\geq\frac{1}{\alpha},$

### v2

@@ Abstract
E-processes enable hypothesis testing with ongoing data collection while maintaining Type I error control. However, when testing multiple hypotheses simultaneously, current $e$-value based multiple testing methods such as e-BH are not invariant to the order in which data are gathered for the different $e$-processes. This can lead to undesirable situations, e.g., where a hypothesis rejected at time $t$ is no longer rejected at time $t+1$ after choosing to gather more data for one or more $e$-processes unrelated to that hypothesis. We argue that multiple testing methods should always work with suprema of $e$-processes. We provide an example to illustrate that e-BH does not control the FDR, at level $\alpha$ when applied to suprema of $e$-processes. From the same example we see that the FWER is not controlled with averaging, and also closed e-BH does not control the FDR. We show that adjusters can be used to ensure FDR-sup control with e-BH under arbitrary dependence.
@@ 1 Introduction
Unfortunately, data gathering under the guarantee (2) is not carefree. Let us illustrate this with an example. For $K=2$ $e$-values, e-BH rejects any hypothesis for which the $e$-value exceeds $2/\alpha$, and both hypotheses if both $e$-values exceed $1/\alpha$. Now suppose that at time $t$ we have $1/\alpha\leq E_{t}^{1}<2/\alpha$ and $E_{t}^{2}<1/\alpha$. Gathering more data for both $e$-processes, however, the situation could reverse, and $E_{t+1}^{1}<1/\alpha$ and $1/\alpha\leq E_{t+1}^{2}<2/\alpha$. Now, the researcher regrets gathering more data for $E^{1}$, since (2) does not support any rejection now, but would have supported rejection of both null hypotheses if the researcher would have stopped $E^{1}$, taking $E^{1}_{t+1}=E^{1}_{t}$. The lack of the carefree property complicates experimental design because collecting more data for one or more hypotheses can cause previously rejected hypotheses—those for which more data has been gathered, or even others—to lose their rejected status. This means that researchers may wish to consider carefully, in real time, for which hypotheses to gather more data, adding data where it seems most promising. As a result, the process becomes volatile and harder to manage. Carefree procedures avoid this problem by ensuring that collecting more data never leads to losing discoveries.
For carefree data gathering, the researcher must always be allowed to revert any $e$-process, for which they regret gathering more data, back to an earlier time point. The optimal reversion is to the maximum so far, so a carefree procedure should guarantee
@@ 2.1 The e-BH procedure
Formally, a non-negative process $E$ is called an $e$-process when $\mathbb{E}[E_{\tau}]\leq 1$ for any stopping time $\tau$, where the expectation is taken under any distribution in the null hypothesis.
We consider the setting in which all $e$-processes are adapted to a common filtration $(\mathcal{F}_{t})_{t\geq 0}$, as this is required to guarantee joint validity under stopping. This setting corresponds to many practical situations, and has been discussed in detail by Wang et al. [2025].
For each $k\in\mathcal{K}$, let $H_{k}$ be associated with an e-process $(E^{k}_{t})_{t\geq 0}$, and let $\tau_{1},\tau_{2},\ldots,\tau_{K}$ be their arbitrary stopping times. Let $S_{k}=E^{k}_{\tau_{k}}$ be the stopped $e$-processes.
@@
The stopped e-BH procedure applied to arbitrarily dependent e-processes, adapted to a common filtration), has FDR at most $K_{0}\alpha/K$.
@@ 2.3 Related work
Online multiple testing has been considered both classically with $p$-values [Fischer et al., 2024] and $e$-values [Fischer and Ramdas, 2024a, b, Xu and Ramdas, 2023], but online is here to be understood in the dimension of the hypothesis space: time points represent the addition of an hypothesis. Until recent, the only paper that considered multiple testing with $e$-values where data can be added sequentially, was [Xu and Ramdas, 2023], where they introduce the doubly sequential framework. For $p$-values there is work on interim-analyses and group-sequential multiple testing, e.g. [Sarkar et al., 2019] on FDR; but to the best of our knowledge there is no work on any-time valid methods. Between uploading the present paper to arxiv and the submission, two notable developments appeared. Wang et al. [2025] provide a rigorous setup for using e-BH on stopped e-processes. They raise the important issue that one needs to be careful about the filtrations used for the stopped e-processes. In the present paper we circumvented the issue by assuming a common underlying filtration, but it would be good to investigate the interplay between their result and ours. Secondly, two papers appeared in the same week: [Goeman et al., 2025] and [Xu et al., 2025]; proposing a novel necessary and sufficient principle for FDR controlling methods based on e-values. The authors have teamed up and are now writing a joint paper from the two preprints. Their principle gives rise to a substantial improvement of e-BH, we call it here closed e-BH, and we added a subsection to show that our central proposition holds for closed e-BH as well. Lastly, Hartog and Lei [2025] study closed testing for familywise error rate (FWER) control, both in the static and sequential setting.
3 Multiple testing with the running maximum of an e-process
@@ 3.1 Proof of Proposition 1
$M_{t}^{1}\geq\frac{2}{\alpha},\quad M_{t}^{2}\geq\frac{2}{\alpha},\quad\text{or both }M_{t}^{1},M_{t}^{2}\geq\frac{1}{\alpha},$ (4)
@@
3.2 Closed e-BH
Recently and in the same week, both Goeman et al. [2025] and Xu et al. [2025] proposed a necessary and sufficient principle for FDR controlling methods. The principle gives rise to an improvement over standard e-BH, which we will now call closed e-BH. In Theorem 2 of Goeman et al. [2025], it is proven that closed e-BH uniformly improves upon standard e-BH, and in particular, that means that the set of rejected hypotheses by e-BH is always included in the sets rejected by closed e-BH. Therefore, the proof of Proposition 1 directly extends to closed e-BH.
Corollary 1.
The procedure of applying closed e-BH to the running maxima of arbitrarily dependent e-processes does not control the FDR at level $\alpha$.
3.3 FWER
Recently, Wang [2024] proved that the only admissible way of merging arbitrary e-values is to use a weighted arithmetic average, which implies that for FWER control under arbitrary dependence, we should average the e-values; mixing with the trivial e-value of $1$ if desired. We see from the events for which one or both hypotheses are rejected in e-BH, stated Equation 4, that averaging those e-values will also always lead to rejection. Since e-values are always positive, if either one of them is larger than $2/\alpha$, the average will exceed $1/\alpha$, and if both of them are larger than $1\alpha$ the average is trivially larger than $1/\alpha$ as well. Therefore, Proposition 1 extends to FWER as well.
Corollary 2.
The procedure of averaging the running maxima of arbitrarily dependent e-processes does not control the FWER at level $\alpha$.
@@ 5 Conclusion
Acknowlegdgemenets
We thank Aaditya Ramdas and Nick Koning for helpful comments on an earlier version of this paper.